OpenStudy (anonymous):

Evaluate -2-5^3-1 ___________ 2^3-2^2

OpenStudy (anonymous):

Does the line represent division?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

5^3= 125 2^3= 8 2^2= 4

OpenStudy (anonymous):

5^3=5*5*5=125 2^3=2*2*2=8 2^2=2*2=4

OpenStudy (anonymous):

Then you wanna plug those in, should get -32.